A ring and a disc of same roll without slipping along a horizontal surface with same velocity if th ekinetic energy of ring is a joules then that of disc is

Grade:11

1 Answers

Arun
5 years ago
This is the solution if the disc and sphere are of same radius.
 
​KE of ring = 1/2Iw^2 + 1/2mv^2 = 1/2[(mR^2)v^2]/R^2 + 1/2mv^2 = 1/2mv^2 + 1/2mv^2 = mv^2 = 8j

KE of disc = 1/2Iw^2 +1/2mv^2 = 1/4[(mR^2)v^2]/R^2 + 1/2mv^2 = 1/4mv^2 + 1/2mv^2 = 3/4mv^2 = 3/4 x 8j = 6j